determin the equation of the line perpendicular to the line 3y+4x=1 passing through the point (-1,2)
Answer by mananth(16946)

4 x + 3 y = 1
Find the slope of this line
3y = -4 x + 1
Divide by 3
y = -4/3 x + 0.33
Compare this equation with y=mx+b
slope m = -4/3
The slope of a line perpendicular to the above line will be the negative reciprocal = 3/4
m1*m2=-1
The slope of the required line will be 0.75 0.75 = 0.75

m= 3/4 ,point (-1,2)
Find b by plugging the values of m & the point in
y=mx+b
2=-0.75 + b
b= 2.75= 9/4
m= 3/4
The required equation is y= 3/ 4 x+9/4